OpenStudy (anonymous):

Need help please If the average family net worth was 240,000, and there were about 7.3 x 10^7 families. Whats the total family net worth

OpenStudy (anonymous):

\[2.4\times 10^5\times 7.3\times 10^7=2.4\times 7.3\times 10^{5+7}\] is a start

OpenStudy (anonymous):

ok thanks so next I will end up wit 17.52 x 10\[17.52\times10^{12}\]

OpenStudy (anonymous):

would the answer be 1.752^13

OpenStudy (anonymous):

yes i think so let me check

OpenStudy (anonymous):

\[1.752\times 10^{13}\] is correct